# Data Management For Large Language Models: A Survey

### v2

@@ 1 Introduction
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=13em,font=,, where level=2text width=14em,font=,, where level=3text width=11em,font=,, [ Data Management, ver [ Pretraining (§2), ver [ Data Quantity (§2.1) [ Scaling Laws [ Kaplan et al. (2020), DeepMind Chinchilla Scaling Law Hoffmann et al. (2022), leaf, text width=33em ] ] [ Data Repetition [ Villalobos et al. (2022), Muennighoff et al. (2023), Hernandez et al. (2022),
Xue et al. (2023), D4 Tirumala et al. (2023), leaf, text width=33em ] ] ] [ Data Quality (§2.2) [ Deduplication [ Lee et al. (2021), Kandpal et al. (2022), Silcock et al. (2022),
@@
SemDeDup Abbas et al. (2023), leaf, text width=33em ] ] [ Quality Filtering [ Gao (2021), Kreutzer et al. (2022), Gunasekar et al. (2023), Li et al. (2023b),
RefinedWeb Penedo et al. (2023), Marion et al. (2023), Longpre et al. (2023b),
@@
MiniPileKaddour (2023), leaf, text width=33em ] ] [ Toxicity Filtering [ Luccioni and Viviano (2021), Xu et al. (2021), Welbl et al. (2021),
@@
Longpre et al. (2023b), leaf, text width=33em ] ] [ Social Bias [ Dodge et al. (2021), Meade et al. (2022), Gururangan et al. (2022)
Feng et al. (2023), leaf, text width=33em ] ] [ Diversity & Age [ Lee et al. (2023a), D2 Pruning Maharana et al. (2023), Longpre et al. (2023b), leaf, text width=33em ] ] ] [ Domain Composition (§2.3) [ Longpre et al. (2023b), CodeGen2 Nijkamp et al. (2023),
SlimPajama-DC Shen et al. (2023), DSIR Xie et al. (2023b),
DoReMi Xie et al. (2023a), DoGE Fan et al. (2023), leaf, text width=33em ] ] [ Data Management Systems (§2.4) [ Data-Juicer Chen et al. (2023a), Oasis Zhou et al. (2023c), leaf, text width=33em ] ] ] [ Supervised Fine-Tuning (§3), ver [ Data Quantity (§3.1) [ Ji et al. (2023), LIMA Zhou et al. (2023a), Yuan et al. (2023),
Chen et al. (2023b), DMT Dong et al. (2023), Song et al. (2023), leaf, text width=33em ] ] [ Data Quality (§3.2) [ Instruction Quality [ INSTRUCTEVAL Chia et al. (2023), LIMA Zhou et al. (2023a),
Ding et al. (2023), Wang et al. (2023d), Li et al. (2023a),
Instruction Mining Cao et al. (2023), Self-Refine Madaan et al. (2023),
SELF Lu et al. (2023a), SelFee Ye et al. (2023), leaf, text width=33em ] ] [ Instruction Diversity [ UltraChat Ding et al. (2023), LIMA Zhou et al. (2023a),
Alpaca Taori et al. (2023), #InsTag Lu et al. (2023b),
Self-Instruct Wang et al. (2023c), Explore-Instruct Wan et al. (2023), leaf, text width=33em ] ] [ Instruction Complexity [ #InsTag Lu et al. (2023b), WizardLM Xu et al. (2023a),
WizardCoder Luo et al. (2023), Orca Mukherjee et al. (2023),
Tree-Instruct Zhao et al. (2023b), CELLO He et al. (2023),
FollowBench Jiang et al. (2023), leaf, text width=33em ] ] [ Prompt Design [ Mishra et al. (2022), Khashabi et al. (2022), Gonen et al. (2022),
Yin et al. (2023b), Kung and Peng (2023), UIT Liang et al. (2023),
@@
Weber et al. (2023), Gudibande et al. (2023), Song et al. (2023), leaf, text width=33em ] ] ] [ Task Composition (§3.3) [ Wei et al. (2021), Wang et al. (2022), Sanh et al. (2022), Chung et al. (2022),
Flan 2022 Longpre et al. (2023a), ELM Jang et al. (2023), Chen et al. (2023b)
DMT Dong et al. (2023), OPT-IML Iyer et al. (2022), Tulu Wang et al. (2023b), leaf, text width=34em ] ] [ Data-Efficient Learning (§3.4) [ AlShikh et al. (2023), Attendu and Corbeil (2023), Ivison et al. (2023),
Instrucion Mining Cao et al. (2023), AlpaGasus Chen et al. (2023c),
OpenChat Wang et al. (2023a), DiverseEvol Wu et al. (2023),
Dynosaur Yin et al. (2023a), MAmmoTH Yue et al. (2023),
DMT Dong et al. (2023), LoBaSS Zhou et al. (2023b),
@@
Xu et al. (2023b), Data-Juicer Chen et al. (2023a), leaf, text width=34em ] ] ] ]

### v3

@@ Appendix E Taxonomy
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=13em,font=,, where level=2text width=14em,font=,, where level=3text width=11em,font=,, [ Data Management, ver [ Pretraining (§2), ver [ Domain Composition (§2.1) [ Longpre et al. (2023b), Nijkamp et al. (2023), Shen et al. (2023),
Xie et al. (2023b), Xie et al. (2023a), Fan et al. (2023), Ye et al. (2024),
Xia et al. (2023), leaf, text width=33em ] ] [ Data Quantity (§2.2) [ Scaling Laws [ Kaplan et al. (2020), Hoffmann et al. (2022), Su et al. (2024), leaf, text width=33em ] ] [ Data Repetition [ Villalobos et al. (2022), Muennighoff et al. (2023), Hernandez et al. (2022),
@@
Xue et al. (2023), Tirumala et al. (2023), leaf, text width=33em ] ] ] [ Data Quality (§2.3) [ Quality Filtering [ Gao (2021), Kreutzer et al. (2022), Gunasekar et al. (2023), Li et al. (2023d),
Penedo et al. (2023), Marion et al. (2023), Longpre et al. (2023b),
Kaddour (2023), Javaheripi and Bubeck (2023), Gan et al. (2023),
Wettig et al. (2024), leaf, text width=33em ] ] [ Deduplication [ Lee et al. (2021), Kandpal et al. (2022), Silcock et al. (2022),
@@
Abbas et al. (2023), leaf, text width=33em ] ] [ Toxicity Filtering [ Luccioni and Viviano (2021), Xu et al. (2021), Welbl et al. (2021),
@@
Longpre et al. (2023b), leaf, text width=33em ] ] [ Diversity & Age [ Lee et al. (2023a), Maharana et al. (2023), Longpre et al. (2023b), leaf, text width=33em ] ] [ Social Bias* [ Dodge et al. (2021), Meade et al. (2022), Gururangan et al. (2022)
Feng et al. (2023), leaf, text width=33em ] ] [ Hallucinations* [ Li et al. (2022a), McKenna et al. (2023), Dziri et al. (2022), leaf, text width=33em ] ] ] [ Relations Among
Different Aspects (§2.4) [ Ge et al. (2024a), Goyal et al. (2024), Bi et al. (2024), Shen et al. (2023),
@@
Longpre et al. (2023b), leaf, text width=33em ] ] ] [ Supervised Fine-Tuning (§3), ver [ Task Composition (§3.1) [ Wei et al. (2021), Wang et al. (2022), Sanh et al. (2022), Chung et al. (2022),
Longpre et al. (2023a), Jang et al. (2023), Chen et al. (2023b), Xia et al. (2024)
Dong et al. (2023), Iyer et al. (2022), Wang et al. (2023b), Ivison et al. (2023),
@@
Lee et al. (2024), leaf, text width=33em ] ] [ Data Quality (§3.2) [ Instruction Quality [ Chia et al. (2023), Zhou et al. (2023a), Li et al. (2023b), Li et al. (2024b)
Ding et al. (2023), Wang et al. (2023d), Li et al. (2023c), Zhou et al. (2023b),
Cao et al. (2023), Madaan et al. (2023), Du et al. (2023), Li et al. (2024c)
Lu et al. (2023a), Ye et al. (2023), Chen et al. (2023c), Li et al. (2023a),
Li et al. (2023e), Bhatt et al. (2024), Chen and Mueller (2024),
Yang et al. (2024), Mekala et al. (2024), He et al. (2024b), Liu et al. (2024), leaf, text width=33em ] ] [ Instruction Diversity [ Ding et al. (2023), Zhou et al. (2023a), Bukharin and Zhao (2023)
Taori et al. (2023), Lu et al. (2023b), Wang et al. (2023c),
Wan et al. (2023), Wu et al. (2023), Ge et al. (2024b), Huang et al. (2024), leaf, text width=33em ] ] [ Instruction Complexity [ Lu et al. (2023b), Xu et al. (2023a), Luo et al. (2023b), Mukherjee et al. (2023),
Zhao et al. (2023b), He et al. (2023), Jiang et al. (2023b), Sun et al. (2024a)
He et al. (2024a), leaf, text width=33em ] ] [ Prompt Design* [ Mishra et al. (2022), Khashabi et al. (2022), Gonen et al. (2022),
Yin et al. (2023b), Kung and Peng (2023), Liang et al. (2023),
Weber et al. (2023), Gudibande et al. (2023), Song et al. (2023), leaf, text width=33em ] ] [ Hallucinations* [ Zhou et al. (2023a), Chen et al. (2023c), Cao et al. (2023), Lee et al. (2023b), leaf, text width=33em ] ] ] [ Data Quantity (§3.3) [ Ji et al. (2023), Zhou et al. (2023a), Yuan et al. (2023), Zhang et al. (2024)
Chen et al. (2023b), Dong et al. (2023), Song et al. (2023), leaf, text width=33em ] ] [ Dynamic
Data-Efficient Learning (§3.4) [ Training Affects Data [ Attendu and Corbeil (2023), AlShikh et al. (2023), Kung et al. (2023), leaf, text width=33em ] ] [ Data Affects Training [ Yin et al. (2023a), Wang et al. (2023a), Dong et al. (2023),
Xu et al. (2023b), Wang et al. (2024), Sun et al. (2024a), Kim and Lee (2024), leaf, text width=33em ] ] ] [ Relations Among
Different Aspects (§2.4) [ Lu et al. (2023b), Bukharin and Zhao (2023), Xu et al. (2023c), leaf, text width=33em ] ] ] ]